the number of fish caught varies directly with the nimber of fisherman. if 18 people can catch 30 fish, then 24 will catch how many fish?
Answer by stanbon(75887)

the number of fish caught varies directly with the nimber of fisherman. if 18 people can catch 30 fish, then 24 will catch how many fish?
-------------
x/24 = 18/30
x = 24(18/30)
x = 4*18/5 = 72/5 = 14 2/5
Rounding down x = 14
===========================
Cheers,
Stan H.